The sound intensity level of a racing car at a distance of 10 m is 130 dB. Assume that the racing car like a small point source and ignore the effects of absorption and reflection. (a) What is the power of the sound emitted by the car? (b) What is the sound intensity at a distance of 10 m from the car? (c) What is the sound intensity due to the car at a distance of 1.0 km from the car? (d) At what distance from the car would the sound intensity fall to the threshold of hearing

1
Expert's answer
2021-12-12T16:40:44-0500

(a) "P=10^{130/10}P_0=10\text{ W/m}^2."

(b) 130 dB by the condition

(c) "L=130-20\log(1000/10)=90\text{ dB}."